The amount of gold jewelry and other products is measured in karats (K), where 24K represents pure gold. The mark 14K on a chain indicates that the ratio between the mass of the gold in the chain and the mass of the chain is 14:24. If a gold ring is marked 18K and it weighs 0.4 oz, what is the value of the gold in the ring if pure gold is valued at $300 per oz?
Answer by stanbon(75887)

The amount of gold jewelry and other products is measured in karats (K), where 24K represents pure gold.
The mark 14K on a chain indicates that the ratio between the mass of the gold in the chain and the mass of the chain is 14:24.
------------------------------------
If a gold ring is marked 18K and it weighs 0.4 oz, what is the value of the gold in the ring if pure gold is valued at $300 per oz?
----------------------
Proportion:
18/24 = x/0.4
x = 0.4(18/24)
x = 0.4(3/4)
x = (3/4)(4/10) = 3/10 = 0.3 oz (this is the amount of gold in the ring)
====================================================================
Cheers,
Stan H.